# Toward Errorless Training ImageNet-1k

### v1

@@ 1. Introduction
With the utilization of GPUs, many advancements in model architecture have been made that greatly increase the accuracy of machine learning models, e.g., [15, 3, 11], and many others. Still, the approach has largely remained the same: do some versions of stochastic gradient descent until the model reaches a sufficiently accurate benchmark ([17, 13, 14, 10]). We believe that our approach presented in this report should provide us with a novel training alternative in the fields of ML and AI.
@@ 5. Results
Training Evaluation. In the GitHub repository [6], in the ImageNet-1k/Data_Transformation directory, the code, Run_Modularize_Data.m reorganizes the dataset into $k=40$ modules, and then partitions each module into $r=2$ groups of training images. For the training accuracy rate, the code, Run_Training_Evaluation_Model.m, evaluates the accuracy for the SGD-type and the GDT-type FNNs, $N_{i,j,s}$, with $1\leq i\leq 17,1\leq j\leq 40,1\leq s\leq 2$, on the modularized training images $B_{i,j,s}$, This is done by first transforming the original $64\times 64$ color images into the input vectors of the first layer of the FNN, and then obtaining the predicted labels of the FNN.
@@
Model Evaluation. As discussed above, when an image is fed to the model, the model must run it through all featured FNNs against the label classes 0 through 999 before making its prediction by the majority-vote protocol. For a single or a small batch of images, this is done by the code, Ru_Evaluation_Model_Individual.m, in the Model_Evaluation folder. But, for the entire dataset of 1.2 million plus images, this program is too slow. Instead, a parallelized code, Run_Evaluation_main_Model.m, is used to run on batches of the dataset to speed up the process. It outputs the predicted labels of the dataset for the feature-defined Model-1 through Model-6, introduced in the previous section. After the predictions are saved, we can run two more files. One file, Run_Evaluation_post_Model.m, finds the accuracy rates of all six models, and the Top-1 accuracy rate, which is defined to be the rate by which a correct prediction is made by the super majority of the votes, namely the first choice of the predictions. The other file, Run_Evaluation_Confusion_Matrix.m, builds the confusion matrices for the models, from which we can find the labels whose entire image classes are classified perfectly. All these tasks for all featured models, derived from the two types of proto-model can be accomplished by the code, Run_Model_Analysis.m in the ImageNet-k folder. The results are shown in Tab.3. It shows the following: The more features a model has, the more accurate it becomes, with a modest drop in the Top-1 rate. Models with a single hidden layer perform better than models with two hidden layers. The Top-1 rate for the GDT-trained featured models is always better than the SGD-trained featured models. For the conjointly trained models for which GDT training is directly based on SGD trained models, featured models of the former always perform better than models of the latter type, as shown by the h2-models. However, when this dependence is removed, i.e., the SGD proto-models are trained beyond the 98% threshold stoppage, their offspring featured models can outperform the GDT kind in model accuracy, but not in the Top-1 accuracy.
@@ 6. Discussion
If double labeling is detected in a dataset, we should be extra cautious when deploying a model trained on it. This is because the problem is bound to cause classification errors, which can cascade down to applications. For this reason, we did not pursue a model evaluation on the validation data since our search algorithm is not equipped to isolate the causes of the prediction error. The dataset’s probable curating errors can cause needless inaccuracy for the GDT-trained models, which can be amplified by the curating errors when the model is applied to the validation dataset. As a result, we used only the simpler kinds of features, consisting of linear and reversible combinations of the primary color channels, to test the GDT method on the dataset.
For image datasets free of double labeling, our GDT method can train models error-free, a vital goal in some applications, especially in the fields of Medicine and Healthcare, where repeating errors must not be tolerated. For problematic image datasets, our results suggest that it may not be worthwhile to add GDT training to SGD training, because the overall performance of our SGD models is on par with the GDT models, and GDT training is considerably more time-consuming than SGD training. We note that our SGD models with six features (Model-6), 113,798,880 parameters, and 566,880 neurons, have already outperformed all conventional SGD methods in the literature.
For practical applications and image datasets free of double labeling, we envision the inclusion of more sophisticated features, such as various convolutions of, or singular value decompositions of images. Such feature inclusions are for future studies to quantify the predictive capability of trained models in terms of the size of neighborhoods of training data in which models’ predictions remain true. Such studies should have a more concise explanation for models’ performance on validation and test data, as well as the true scope of models’ capabilities and limitations. In addition, only for consistent datasets can we pursue the smallest FNN model that can be trained fully. For the MINST handwritten digits, we only need one hidden layer of 20 neurons and 15,910 parameters to achieve that goal ([5]). We believe that the size of our model for ImageNet-1k can be reduced by one order of magnitude for the same level of performance, which we did not pursue in this study.
@@
Because our model is structured in parallel FNNs with minimal hidden layers and modest batch sizes, training is limited only by the number of CPU cores. In the ideal setting where each FNN is assigned to a core, an SGD proto-model can be trained in 1 hour, and a GDT proto-model can be trained in 3 hours, which is the same length of time to update any FNN. As a demonstration, no GPUs were used to train our model.
@@
Data Availability: All model weights and evaluation codes can be found in the GitHub repository, [6]. For questions about the training methods, send emails to the first author.

### v4

@@ 1. Introduction
With the utilization of GPUs, many advancements in model architecture have been made that greatly increase the accuracy of machine learning models, e.g., [15, 11], and many others. Still, the approach has largely remained the same: do some versions of stochastic gradient descent until the model reaches a sufficiently accurate benchmark ([17, 13, 14, 10]). We believe that our approach presented in this report should provide us with a novel training alternative in the fields of ML and AI.
@@ 5. Results
Training Evaluation. In the Hugging Face repository [6], in the folder of Data_Transformationfig, the code, Run_Modularize_Data.m reorganizes the dataset into $k=40$ modules, and then partitions each module into $r=2$ groups of training images. For the training accuracy rate, the code, Run_Training_Evaluation_Model.m, evaluates the accuracy for the SGD-type and the GDT-type FNNs, $N_{i,j,s}$, with $1\leq i\leq 17,1\leq j\leq 40,1\leq s\leq 2$, on the modularized training images $B_{i,j,s}$, This is done by first tranfsforming the original $64\times 64$ color images into the input vectors of the first layer of the FNN, and then obtaining the predicted labels of the FNN.
@@
Model Evaluation. As discussed above, when an image is fed to the model, the model must run it through all featured FNNs against the label classes 0 through 999 before making its prediction by the majority-vote protocol. For a single or a small batch of images, this is done by the code, Ru_Evaluation_Model_Individual.m, in the Model_Evaluation folder. But, for the entire dataset of 1.2 million plus images, this program is too slow. Instead, a parallelized code, Run_Evaluation_main_Model.m, is used to run on batches of the dataset to speed up the process. It outputs the predicted labels of the dataset for the feature-defined Model-1 through Model-6, introduced in the previous section. After the predictions are saved, we can run two more files. One file, Run_Evaluation_post_Model.m, finds the accuracy rates of all six models, and the Top-1 accuracy rate, which is defined to be the rate by which a correct prediction is made by the super majority of the votes, namely the first choice of the predictions. The other file, Run_Evaluation_Confusion_Matrix.m, builds the confusion matrices for the models, from which we can find the labels whose entire image classes are classified perfectly. All these tasks for all featured models, derived from the two types of proto-model can be accomplished by the code, Run_Model_Analysis.m in the ImageNet-1k folder. The results are shown in Tab.3. It shows the following: The more features a model has, the more accurate it becomes, with a modest drop in the Top-1 rate. Models with a single hidden layer perform better than models with two hidden layers. The Top-1 rate for the GDT-trained featured models is always better than the SGD-trained featured models. For the conjointly trained models for which GDT training is directly based on SGD-trained models, featured models of the former always perform better than models of the latter type, as shown by the h2-models. However, when this dependence is removed, i.e., the SGD proto-models are trained beyond the 98% threshold stoppage, their offspring featured models can outperform the GDT kind in model accuracy, but not in the Top-1 accuracy.
@@
featured T_h1_m1 T_h1_m2 T_h2_m1
model Acc. Top-1 Acc. Top-1 Acc. Top-1
Model-1 74.843 100 72.427 100 41.807 100
Model-2 94.755 100 94.456 100 84.018 100
Model-3 97.012 100 96.921 100 92.438 99.997
Model-4 97.714 100 97.728 100 94.807 99.992
Model-5 98.026 100 98.022 99.998 95.939 99.987
Model-6 98.213 99.999 98.224 99.999 96.537 99.983
mean 93.427 99.999 92.963 99.999 84.258 99.993
S_h1_m1 S_h1_m2 S_h2_m1
Model-1 80.573 100 80.435 100 45.154 100
Model-2 95.608 99.995 95.743 99.997 81.292 99.991
Model-3 96.952 99.949 97.093 99.960 89.632 99.925
Model-4 97.383 99.851 97.570 99.876 92.371 99.804
Model-5 97.549 99.741 97.748 99.757 93.758 99.674
Model-6 97.669 99.635 97.866 99.663 94.569 99.562
mean 94.289 99.862 94.409 99.876 82.796 99.826
Table 5. Accuracy rates for the expanded majority voting protocol.
Specifically, the majority voting protocol for each featured model consists of three voting steps. Let $p$ be the number of features for the model. For Model-1, $p=3$; for Model-2, $p=6$, and so on. First, for each input image, we choose the top label output from every FNN $N_{i,j,s}$ of the featured model, where $i$ ranges from 1 to the model defining $p$ features. In the second step, the selection is carried out by the module-submodule, $n_{j,s}$, over the set of $p$ features, resulting in $kr=80$ selected labels together with the number of votes each winning label gets. In the ideal situation where all FNNs are perfectly trained, the module-submodule that is assigned to train the image will return the true label together with the super-majority vote number $p$. In contrast, for any other module-submodule, its selected label’s winning number is between 1 and $p$. It is possible that such a wrong label can have the super-majority winning number $p$. In the third step, selection is carried out for the set of winning numbers. Again, for the ideal situation, if no other module-submodules reach the super-majority winning number $p$, the final output will be the true label by the module-submodule that is trained on the image. Otherwise, if there is a tie because a wrong label also has the super-majority number $p$ by chance, we first compute the variance of the loss function values over the features for each of the tying module-submodules, and then choose the module-submodule having the smallest variance to break the tie. Table 3 is produced by this procedure.
Table 5 is produced the same way except for the following modification for the first step. Rather than choosing only the top label output from every FNN $N_{i,j,s}$ of the featured model, we use instead the top three labels from every FNN, with the effect of increasing the pool of candidate labels for the first round of selection. This is because the intended label may not always be the top choice of a given feature, but is more likely among the top two or top three choices. That is, the bigger the pool, the more likely the true label will reach the super-majority. The results show that the Top-1 rate improves considerably with a minor drop in the overall accuracy. For example, for Model-5 of T_h1_m1, every correctly predicted label is by the intended super-majority, namely, the winning number of 15 votes. If the dataset is consistent without double labeling, the accuracy rate should be significantly higher, possibly to 100%, the same as the Top-1 rate.
@@ 6. Discussion
If double labeling is detected in a dataset, we should be extra cautious when deploying a model trained on it. This is because the problem is bound to cause classification errors, which can cascade down to applications. Specifically, for datasets free of contractions, the accuracy rate for GDT-trained models in Tab.2 should be 100%. Otherwise, the accuracy statistics for featured models in Tab.3 will be muddled up. Due to these reasons, we did not pursue a model evaluation on the validation data since our search algorithm is not equipped to isolate the causes of the prediction error. The dataset’s probable curating errors can cause needless inaccuracy for the GDT-trained models, which will ripple through the featured models. As a result, we used only the simpler kinds of features, consisting of linear and reversible combinations of the primary color channels, to test the GDT method on the dataset.
For image datasets free of double labeling, our GDT method can train models error-free, a vital goal in some applications, especially in the fields of Medicine and Healthcare, where repeating errors must not be tolerated. For problematic image datasets, our results suggest that it may not be worthwhile to add GDT training to SGD training, because the overall performance of our SGD models is on par with the GDT models, and GDT training is considerably more time-consuming than SGD training. We note that our SGD models with six features (Model-2), 113,798,880 parameters, and 566,880 neurons, have already outperformed all conventional SGD methods in the literature. For practical applications and image datasets free of double labeling, we envision the inclusion of more sophisticated features, such as various convolutions of, or singular value decompositions of images. Such feature inclusions are for future studies to quantify the predictive capability of trained models in terms of the size of neighborhoods of training data in which models’ predictions remain true. Such studies should have a more concise explanation for models’ performance on validation and test data, as well as the true scope of models’ capabilities and limitations. In addition, only for consistent datasets can we pursue the smallest FNN model that can be trained fully. For the MINST handwritten digits, we only need one hidden layer of 20 neurons and 15,910 parameters to achieve that goal ([4]). We believe that the size of our model for ImageNet-1k can be reduced by one order of magnitude for the same level of performance, which we did not pursue in this study.
@@
Because our model is structured in parallel FNNs with minimal hidden layers and modest batch sizes, training is limited only by the number of CPU cores. In the ideal setting where each FNN is assigned to a core, an SGD proto-model can be trained in 1 hour, and a GDT proto-model can be trained in 3 hours, which is the same length of time to update any FNN. As a demonstration, no GPUs were used to train our model. However, we expect that our current training speed can be considerably boosted by incorporating GPU parallelism into our training code.
@@
Data Availability: All model weights and evaluation codes can be found in the Hugging Face repository, [6]. For questions about the training methods, send emails to the first author.